OpenStudy (anonymous):

CaN YoU SolVE IT : Lawrence takes 8 subjects at school, and for each one he receives a code from 1 to 7. If the sum of the codes is fifty then the least no. of 7s he must have received is: A) 6 B) 5 C) 4 D) 3 E) 2

OpenStudy (anonymous):

6*8=48

OpenStudy (anonymous):

So he must get at least two 7

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Okay now !! @sauravshakya u are wrong

OpenStudy (anonymous):

Didn't I say two 7

OpenStudy (anonymous):

Option E

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

OKAY!! its l'll tuff or easy for u

OpenStudy (anonymous):

If he didnt get any 7 then the highest he can get is 6 in all subject. So, 6*8=48 But the sum is 50.. So, 2 must be added on both sides 6*8+2=48+2 6*6+7*2=50

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Good to see that you were corrreeeeccttt!!! @sauravshakya

OpenStudy (anonymous):

Thanx